### v1

@@ 6 A Proof of Theorem˜3
In this section, we give a proof of Theorem˜3 that goes through the topology of hockey-stick curves.
As we have seen, Theorem˜3 can be proved by showing that, for any family $\mathcal{F}$ that is not well-ordered, we can find $g_{1},g_{2}\in\mathcal{F}^{\infty}$ for which
$g_{1}\otimes\operatorname{conv}\{g_{1},g_{2}\}\prec\operatorname{conv}\{g_{1}\otimes g_{1},g_{1}\otimes g_{2}\}$
Equivalently, in the language of PLDs, we would like to show that if $L_{1}$ and $L_{2}$ are not well-ordered, then $L_{1}\oplus\sup\{L_{1},L_{2}\}\succ\sup\{L_{1}\oplus L_{1},L_{1}\oplus L_{2}\}$.
@@
Our first step is to show that gaps between PLDs can be amplified in some sense by composition. For a distribution $A$ in the reals, we denote by $\operatorname{supp}(A):=\{t\in\mathbb{R}:\forall\text{open }U\ni t,A(U)>0\}$ the support of $A$, namely the points with locally positive probability mass. It turns out that any PLD whose support can be shifted to hit gaps between other PLDs causes the gaps to align when composed.
@@
Let $A,B,B_{1},B_{2}$ be PLDs so that $B_{i}\prec B$ with $h_{B_{i}}(e^{\varepsilon_{i}+u})<h_{B}(e^{\varepsilon_{i}+u})$ for some $\varepsilon_{1},\varepsilon_{2}\in\operatorname{supp}(A)$ and some $u\in\mathbb{R}$. Then
$A\oplus B\succ\sup\{A\oplus B_{1},A\oplus B_{2}\}.$
@@
First, it is clear that $A\oplus B\succeq\sup\{A\oplus B_{1},A\oplus B_{2}\}$, so we just need to find a single gap.
To that end, by continuity, we can find a constant $\rho>0$ s.t $\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{B}(e^{\varepsilon})-h_{B_{i}}(e^{\varepsilon})>0$. Now, set $\varepsilon^{\star}:=\varepsilon_{1}+\varepsilon_{2}+u$ and notice that
$h_{A\oplus B}(e^{\varepsilon^{\star}})-h_{A\oplus B_{i}}(e^{\varepsilon^{\star}})$ $=\mathbb{E}_{Z\sim A\oplus B}[(1-e^{\varepsilon^{\star}-Z})_{+}]-\mathbb{E}_{Z\sim A\oplus B_{i}}[(1-e^{\varepsilon^{\star}-Z})_{+}]$
$=\mathbb{E}_{Z_{1}\sim A}[\mathbb{E}_{Z_{2}\sim B}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]-\mathbb{E}_{Z_{2}\sim B_{i}}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]]$
$=\mathbb{E}_{Z\sim A}[h_{B}(e^{\varepsilon^{\star}-Z})-h_{B_{i}}(e^{\varepsilon^{\star}-Z})]$
$\geq\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{2-i}\pm\rho)}h_{B}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})-h_{B_{i}}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})$
$=\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{B}(e^{\varepsilon})-h_{B_{i}}(e^{\varepsilon})$
@@
In particular, $h_{\sup\{A\oplus B_{1},A\oplus B_{2}\}}(e^{\varepsilon^{\star}})=\max\{h_{A\oplus B_{1}}(e^{\varepsilon^{\star}}),h_{A\oplus B_{2}}(e^{\varepsilon^{\star}})\}<h_{A\oplus B}(e^{\varepsilon^{\star}})$. ∎
In order to apply gap amplification to PLDs that are not well-ordered i.e. $L_{1},L_{2}\prec\sup\{L_{1},L_{2}\}$, we would like to reduce the problem to geometric and topological properties of convex curves over $\mathbb{R}^{\times}:=(0,\infty)$. To that end, we also notice that the support of a PLD is determined by the topology of its hockey-stick curve. This result follows immediately from Proposition˜5
@@
Our problem is now purely geometric. Given decreasing, convex, unordered curves $h_{1},h_{2}:\mathbb{R}^{\times}\to[0,1]$ that meet in the limit $x\to 0$, we would like to show that one of them can be scaled along the x-axis so that its elbows hit a region where $h_{1}<h_{2}$ and another region where $h_{1}>h_{2}$.
@@
Critically, non-trivial closed and divisible subsets of $\mathbb{R}^{\times}$ have a highly regular multiplicative structure enclosed byt their divisors.
@@
We now can prove our main result of the section, which implies Theorem˜3. We note the small caveat that the queries must be non-degenerate, namely their PLDs must be supported at at least two finite points. This is not serious practical limitation.
@@
Let $h_{1},h_{2}:\mathbb{R}^{\times}\to\mathbb{R}$ be non-degenerate hockey-stick curves that are not well-ordered. Then one of $\operatorname{supp}(h_{1}),\operatorname{supp}(h_{2})$ links $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$.
@@
We begin with the easy case: Assume that one of $\partial X(h_{1},h_{2}),\partial X(h_{2},h_{1})$ is not fully contained in $C(h_{1},h_{2})$. Without loss of generality, suppose $\partial X(h_{1},h_{2})\nsubseteq C(h_{1},h_{2})$. By Proposition 17, $\operatorname{supp}(h_{2})$ meets $\partial X(h_{1},h_{2})\subseteq\overline{X(h_{1},h_{2})}$. Furthermore, since $h_{1}$ and $h_{2}$ are not well-ordered, $X(h_{2},h_{1})\neq\varnothing,\mathbb{R}^{\times}$ and thus $\operatorname{supp}(h_{2})$ meets $X(h_{2},h_{1})$ as well by Proposition 18. By Lemma 2, $\operatorname{supp}(h_{2})$ must link $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$.
@@
In particular, every root of $f$ lies in $\mathbb{R}^{\times}$. On the other hand, a homogeneous linear recurrence over the integers admits a bounded non-constant solution if and only its characteristic polynomial has a non-unit root lying on the complex unit circle (see e.g. Section 2.3 of [elaydi2005introduction]). Therefore either $y_{k}=h_{1}(a^{k}b_{1})=h_{2}(a^{k}b_{1})$ is unbounded, which is impossible because hockey-stick curves are bounded, or it must be constant, which is also impossible because hockey-stick curves are monotone, so this would force $h_{1}$ and $h_{2}$ to be constant and thus well-ordered. ∎
Theorem˜3 now follows immediately by combining Propositions 14 and 15 with Theorem 4 by taking $B:=\sup\{B_{1},B_{2}\}$.
Appendix A $f$-DP Results, Proofs and Special Cases
We present additional results focusing on $f$-DP, which we use to instantiate our main results Theorems˜1, 2 and 3 on special cases, that may help with intuition. For completeness, Algorithm˜2 states the privacy filter algorithm for $f$-DP, and is a direct analogy to Algorithm˜1.
Algorithm 2 NaturalFilter ($f$-DP)
0: Privacy bound $f$, adversary (or analyst) $\mathcal{A}$, family of tradeoff curves $\mathcal{F}$, query capacity $k$
$i=0$
while $\mathcal{A}$ continues and $i<k$ do
  $i\leftarrow i+1$
  $\mathcal{A}$ gives mechanism $\mathcal{M}_{i}$ that is $f_{i}$-DP, with $f_{i}\in\mathcal{F}$ $\triangleright$ can depend on previous results $Y_{<i}$
  if $f_{1}\otimes\ldots\otimes f_{i}\geq f$ then
   $\mathcal{A}$ receives $Y_{i}\sim\mathcal{M}_{i}$
  end if
end while
return $(Y_{1},\ldots,Y_{i})$
A.1 Proof of Proposition˜12
Proof.
For each $i\in\{1,2\}$, let $f_{i}$ be a tradeoff curve with $n_{i}$ linear segments and $f_{i}(0)=1-\delta_{i}$. Each linear segment $j\in[n_{i}]$ has a slope $f^{\prime}_{i,j}\leq 0$ over an interval $\alpha_{i,j},\alpha_{i,j+1})$. The width of interval $j$ is denoted by $w_{i,j}$. We also assume $f^{\prime}_{i,1}\geq-f^{\prime}_{i,2}\geq\cdots\geq-f^{\prime}_{i,n_{i}}$. Note that we have $\alpha_{i,1}=0$, $\alpha_{i,n_{i}+1}=1$, $f_{i}(1)=0$, and $\sum_{j=1}^{n_{i}}w_{i,j}=1$.
@@ A.1 Proof of Proposition˜12
A.2 A Simple Derivation of Theorem˜3 for Approximate DP
Definition 8 (Crossing tradeoff curves).
We say that two tradeoff curves $f,g\in\mathcal{T}$ cross if there exists $\alpha_{1},\alpha_{2}\in[0,1]$ s.t.:
$f(\alpha_{1})>g(\alpha_{1})\ \textrm{and}\ f(\alpha_{2})<g(\alpha_{2}).$
By continuity, there has to be $\alpha\in(\alpha_{1},\alpha_{2})$ s.t. $f(\alpha)=g(\alpha)$. For any two $(\varepsilon,\delta)$-DP with tradeoff curves that cross, there cannot be free $f$-DP filters.
We can now prove Proposition˜13, a special case of Theorem˜3 for $(\varepsilon,\delta)$-DP.
@@ A.2 A Simple Derivation of Theorem˜3 for Approximate DP
Consider $g_{1}\triangleq f_{\varepsilon_{1},\delta_{1}}$ and $g_{2}\triangleq f_{\varepsilon_{2},\delta_{2}}$ that cross. By symmetry w.r.t. the $y=x$ axis, they have to cross at an $\alpha<\min\{\alpha_{1}^{\star},\alpha_{2}^{\star}\}$, where $\alpha_{i}^{\star}\ \textrm{s.t.}\ \alpha_{i}^{\star}=g_{i}(\alpha_{i}^{\star})$ (otherwise they would never cross). This means that their linear segments cross: one function starts strictly lower, and then ends-up strictly higher by $\min\{\alpha_{1}^{\star},\alpha_{2}^{\star}\}$ (again, otherwise they would never cross). Without loss of generality, we reorder functions so that $g_{2}$ starts strictly lower.
We now know that $\delta_{2}>\delta_{1}\geq 0$ ($g_{2}$ starts lower); $\varepsilon_{1}>\varepsilon_{2}\geq 0$ ($g_{1}$ has a steeper slope to catch-up to $g_{2}$ so that they can cross); and $\alpha_{2}^{\star}>\alpha_{1}^{\star}$. The last inequality holds because $g_{1}$ is strictly lower by the time is crosses $y=x$ at $\alpha_{1}^{\star}$, so at that point $g_{2}(\alpha_{1}^{\star})>\alpha_{1}^{\star}$, and by continuity of $g_{2}$ we have that $\alpha_{2}^{\star}>\alpha_{1}^{\star}$. Since we are dealing with approximate DP curves, we also know that $\alpha_{i}^{\star}=\frac{1-\delta_{i}}{1+e^{\varepsilon_{i}}}$.
Now, in Theorem˜2, pick $\mathcal{H}=\{g_{1},g_{2}\}\subseteq\mathcal{F}$ and $g=g_{1}\in\mathcal{H}\subseteq\mathcal{F}$. We make a few useful observations about the first segment of $g_{1}\otimes g_{1}$, $g_{1}\otimes g_{2}$, and $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$, by applying Proposition˜12:
The first segment of $g_{1}\otimes g_{1}$ starts at $(1-\delta_{1})^{2}$, with slope $e^{2\varepsilon_{1}}$ for $\alpha\in[0,\alpha_{1}^{\star 2}]$. The second slope is $e^{\varepsilon_{1}}e^{-\varepsilon_{1}}=-1$.
The first segment of $g_{1}\otimes g_{2}$ starts at $(1-\delta_{1})(1-\delta_{2})$, with slope $e^{\varepsilon_{1}+\varepsilon_{2}}$ for $\alpha\in[0,\alpha_{1}^{\star}\alpha_{2}^{\star}]$. The second slope is $e^{\varepsilon_{1}-\varepsilon_{2}}<-1$.
The first segment of $\operatorname{conv}(g_{1},g_{2})$ starts at $(1-\delta_{2})$ (the lowest point at $\alpha=0$), and since $g_{1}$ and $g_{2}$ cross, it meets $g_{1}$ at $\alpha=\alpha_{1}^{\star}$, where its value is $g_{1}(\alpha_{1}^{\star})=1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}$. As a result, the first slope of $\operatorname{conv}(g_{1},g_{2})$ is $\frac{1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}-(1-\delta_{2})}{\alpha_{1}^{\star}}=\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{\varepsilon_{1}}$.
The first segment of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ starts at $(1-\delta_{1})(1-\delta_{2})$, with first slope $\frac{\delta_{1}-\delta_{2}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}$ for $\alpha\in[0,\alpha_{1}^{\star 2}]$.
Let us study the first segment of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ and compare it to that of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$. We first reason about the possible shapes of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$. We saw that $g_{1}\otimes g_{1}(0)=(1-\delta_{1})^{2}>(1-\delta_{1})(1-\delta_{2})=g_{1}\otimes g_{2}(0)$, so $g_{1}\otimes g_{2}$ starts lower. Since $e^{2\varepsilon_{1}}<-e^{\varepsilon_{1}+\varepsilon_{2}}<-e^{\varepsilon_{1}-\varepsilon_{2}}$, $g_{1}\otimes g_{1}<-1$, we know that $g_{1}\otimes g_{1}$ starts with a steepest slope, which stops first (at $\alpha_{1}^{\star 2}<\alpha_{1}^{\star}\alpha_{2}^{\star}$) and then gets the smallest second slope (until at least $\alpha\geq\alpha_{g_{1}\otimes g_{2}}^{\star}$). As a result, either that first slope is steep enough such that $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})<g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and the composed curves cross, or they will never cross and $g_{1}\otimes g_{2}\leq g_{1}\otimes g_{1}$. We treat each case in turn.
Case 1: $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})<g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and the composed curves cross (Figure˜1(a)). In that case, we know that $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha_{1}^{\star 2})=g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})=(1-\delta_{1})^{2}-e^{2\varepsilon_{1}}\alpha_{1}^{\star 2}$. As a result, the slope of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ in $\alpha\in[0,\alpha_{1}^{\star 2}]$ is $\frac{(1-\delta_{1})^{2}-e^{2\varepsilon_{1}}\alpha_{1}^{\star 2}-(1-\delta_{1})(1-\delta_{2})}{\alpha_{1}^{\star 2}}=\frac{(1-\delta_{1})(\delta_{2}-\delta_{1})}{\alpha_{1}^{\star 2}}-e^{2\varepsilon_{1}}=(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{2\varepsilon_{1}}$ (the last step uses $\alpha_{i}^{\star}=\frac{1-\delta_{i}}{1+e^{\varepsilon_{i}}}$).
Remember that $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(0)=(1-\delta_{1})(1-\delta_{2})=\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$, and the first slope of with first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})=\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}<(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{2\varepsilon_{1}}$ since $\delta_{2}-\delta_{1}>0$.
We have shown that in this case, $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})>\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ at least on $(0,\alpha_{1}^{\star 2}]$, so by Theorem˜2, and there are no $f$-DP filters.
Case 2: $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})\geq g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})=g_{1}\otimes g_{2}$ (Figure˜1(b)). Intuitively, we know that the first slope of $\operatorname{conv}(g_{1},g_{2})$ is steeper than that of $g_{2}$, since $g_{2}$ starts lower and crosses with $g_{1}$. Formally, the first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ is $\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}=-e^{\varepsilon_{1}}\big(e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)$. We know that $1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}=g_{1}(\alpha_{1}^{\star})<g_{2}(\alpha_{1}^{\star 2})=1-\delta_{2}-e^{\varepsilon_{2}}\alpha_{1}^{\star}$, so $\delta_{2}-\delta_{1}<\alpha_{1}^{\star}(e^{\varepsilon_{1}}-e^{\varepsilon_{2}})$ and $0<e^{\varepsilon_{2}}<e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}$. That is, $e^{\varepsilon_{1}}\big(e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)<-e^{\varepsilon_{1}}e^{\varepsilon_{2}}$.
In this case as well, the first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ is strictly more steep than that of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$, so $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})>\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ at least on $(0,\alpha_{1}^{\star 2}]$. By Theorem˜2, and there are no free $f$-DP filters. ∎

### v2

@@ 8 A Proof of Lemma˜4
In this section, we give a proof of the key technical Lemma˜4 that goes through the topology of hockey-stick curves. That is, we are given non-degenerate PLDs $L_{1}$ and $L_{2}$ that are not ordered. Our goal is to argue that, for one of these two PLDs $L_{i}$, $L_{i}\oplus\sup\{L_{1},L_{2}\}\succ\sup\{L_{i}\oplus L_{1},L_{i}\oplus L_{2}\}$.
@@
Our first step is to show that gaps between PLDs can be amplified in some sense by composition. Recall that the support of a distribution $A$ over $\overline{\mathbb{R}}$, is $\operatorname{supp}(A):=\{t\in\mathbb{R}:\forall\text{open }U\ni t,A(U)>0\}$, i.e. the finite points with locally positive probability. It turns out that any PLD whose support can be shifted to hit gaps between other PLDs causes the gaps to align when composed.
@@
Let $A,L,L_{1},L_{2}$ be PLDs so that $L_{i}\prec L$ with $h_{L_{i}}(e^{\varepsilon_{i}+u})<h_{L}(e^{\varepsilon_{i}+u})$ for some $\varepsilon_{1},\varepsilon_{2}\in\operatorname{supp}(A)$ and some $u\in\mathbb{R}$. Then
$A\oplus L\succ\sup\{A\oplus L_{1},A\oplus L_{2}\}.$
@@
First, it is clear that $A\oplus L\succeq\sup\{A\oplus L_{1},A\oplus L_{2}\}$, so we just need to find a single gap.
To that end, by continuity, we can find a constant $\rho>0$ s.t $\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{L}(e^{\varepsilon})-h_{L_{i}}(e^{\varepsilon})>0$. Now, set $\varepsilon^{\star}:=\varepsilon_{1}+\varepsilon_{2}+u$ and notice that
$h_{A\oplus L}(e^{\varepsilon^{\star}})-h_{A\oplus L_{i}}(e^{\varepsilon^{\star}})$ $=\mathbb{E}_{Z\sim A\oplus L}[(1-e^{\varepsilon^{\star}-Z})_{+}]-\mathbb{E}_{Z\sim A\oplus L_{i}}[(1-e^{\varepsilon^{\star}-Z})_{+}]$
$=\mathbb{E}_{Z_{1}\sim A}[\mathbb{E}_{Z_{2}\sim L}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]-\mathbb{E}_{Z_{2}\sim L_{i}}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]]$
$=\mathbb{E}_{Z\sim A}[h_{L}(e^{\varepsilon^{\star}-Z})-h_{L_{i}}(e^{\varepsilon^{\star}-Z})]$
$\geq\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{2-i}\pm\rho)}h_{L}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})-h_{L_{i}}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})$
$=\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{L}(e^{\varepsilon})-h_{L_{i}}(e^{\varepsilon})$
@@
In particular, $h_{\sup\{A\oplus L_{1},A\oplus L_{2}\}}(e^{\varepsilon^{\star}})=\max\{h_{A\oplus L_{1}}(e^{\varepsilon^{\star}}),h_{A\oplus L_{2}}(e^{\varepsilon^{\star}})\}<h_{A\oplus L}(e^{\varepsilon^{\star}})$. ∎
In order to apply gap amplification to PLDs that are not totally-ordered i.e. $L_{1},L_{2}\prec\sup\{L_{1},L_{2}\}$, we would like to reduce the problem to geometric and topological properties of convex curves over $\mathbb{R}^{\times}:=(0,\infty)$. To that end, we also notice that the support of a PLD is determined by the topology of its hockey-stick curve. This result follows immediately from Proposition˜5
@@
Our problem is now purely geometric. Given decreasing, convex, unordered curves $h_{1},h_{2}:\mathbb{R}^{\times}\to[0,1]$ that meet in the limit $x\to 0$ (recall Proposition˜4), we show that one of them can be scaled along the x-axis so that its elbows hit a region where $h_{1}<h_{2}$ and another region where $h_{1}>h_{2}$.
@@
Critically, non-trivial closed and divisible subsets of $\mathbb{R}^{\times}$ have a highly regular multiplicative structure enclosed by their divisors.
@@
We now can prove our main result of the section, which implies Lemma˜4. We note the small caveat that the queries must be non-degenerate, namely their PLDs must be supported at at least two finite points. This is not serious practical limitation.
@@
Let $h_{1},h_{2}:\mathbb{R}^{\times}\to\mathbb{R}$ be non-degenerate hockey-stick curves that are not totally-ordered. Then one of $\operatorname{supp}(h_{1}),\operatorname{supp}(h_{2})$ links $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$.
@@
We begin with the easy case: Assume that one of $\partial X(h_{1},h_{2}),\partial X(h_{2},h_{1})$ is not fully contained in $C(h_{1},h_{2})$. Without loss of generality, suppose $\partial X(h_{1},h_{2})\nsubseteq C(h_{1},h_{2})$. By Proposition 16, $\operatorname{supp}(h_{2})$ meets $\partial X(h_{1},h_{2})\subseteq\overline{X(h_{1},h_{2})}$. Furthermore, since $h_{1}$ and $h_{2}$ are not totally-ordered, $X(h_{2},h_{1})\neq\varnothing,\mathbb{R}^{\times}$ and thus $\operatorname{supp}(h_{2})$ meets $X(h_{2},h_{1})$ as well by Proposition 17. By Lemma 13, $\operatorname{supp}(h_{2})$ must link $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$.
@@
In particular, every root of $f$ lies in $\mathbb{R}^{\times}$. On the other hand, a homogeneous linear recurrence over the integers admits a bounded non-constant solution if and only its characteristic polynomial has a non-unit root lying on the complex unit circle (see e.g. Section 2.3 of Elaydi (2005)). Therefore either $y_{k}=h_{1}(a^{k}b_{1})=h_{2}(a^{k}b_{1})$ is unbounded, which is impossible because hockey-stick curves are bounded, or it must be constant, which is also impossible because hockey-stick curves are monotone, so this would force $h_{1}$ and $h_{2}$ to be constant and thus totally-ordered. ∎
Lemma˜4 now follows immediately by combining Propositions 13 and 14 with Theorem 18 by taking $L:=\sup\{L_{1},L_{2}\}$.
Appendix A Proofs from Preliminaries
Proof of Proposition˜5.
$H_{x}(P\parallel Q)$ $=\sup_{E\subseteq\Omega}P(E)-xQ(E)$
$=\sup_{E\subseteq\Omega}\int_{E}1-x/\frac{dP}{dQ}(\omega)\,dP(\omega)$
$=\int_{\Omega}\left(1-x/\frac{dP}{dQ}(\omega)\right)_{+}\,dP(\omega)$
$=\int_{\Omega}(1-xe^{-\log(\ell(\omega))})_{+}\,dP(\omega)$
Proof of Proposition˜6.
By Proposition˜5 and the law of total expectation, it follows that
Proof of Proposition˜9.
First, let $L$ denote the PLD of $(P,Q)$ and let $F$ be its CDF. For $\alpha\in[0,1]$, let $\psi_{\alpha}$ denote the threshold as in Proposition˜8. For any $x\in\mathbb{R}^{\times}$, Proposition˜5 yields
$h(x)$ $=\mathop{\mathbb{E}}_{Z\sim L}[1(Z>\log{x})\cdot(1-xe^{-Z})]$
$=\sup_{\alpha\in[0,1]}\mathop{\mathbb{E}}_{Z\sim L}[\psi_{\alpha}(Z)\cdot(1-xe^{-Z})]$ ($\alpha=F(\log{x})$)
$=\sup_{\alpha\in[0,1]}1-\alpha-x\tau(\alpha)$
$=\sup_{\beta\in[0,1]}1-\tau^{-1}(\beta)-x\beta$ ($\alpha=\tau^{-1}(\beta),\beta=\tau(\alpha)$)
$=1+(\tau^{-1})^{*}(-x).$
Appendix B Proofs for Section 6
Proof of Proposition˜12.
For each $i\in\{1,2\}$, let $f_{i}$ be a tradeoff curve with $n_{i}$ linear segments and $f_{i}(0)=1-\delta_{i}$. Each linear segment $j\in[n_{i}]$ has a slope $f^{\prime}_{i,j}\leq 0$ over an interval $\alpha_{i,j},\alpha_{i,j+1})$. The width of interval $j$ is denoted by $w_{i,j}$. By convexity, we have $f^{\prime}_{i,1}\geq-f^{\prime}_{i,2}\geq\cdots\geq-f^{\prime}_{i,n_{i}}$. Note that we have $\alpha_{i,1}=0$, $\alpha_{i,n_{i}+1}=1$, $f_{i}(1)=0$, and $\sum_{j=1}^{n_{i}}w_{i,j}=1$.
@@ Appendix B Proofs for Section 6
Appendix C On Symmetric PLDs and Dominating Pairs
For the purpose of accurate privacy accounting, we often need to identify the “worst” pairs of datasets and its corresponding privacy loss. This concept can be formalized via the notion of (worst-case) dominating pairs as follows:
Definition 16 (Dominating Pair [58]).
$(P,Q)$ is a dominating pair for a mechanism $\mathcal{M}$ if $(P,Q)\succeq(\mathcal{M}(D),\mathcal{M}(D^{\prime}))$ for all neighbouring datasets $D,D^{\prime}$. Furthermore, it is said to be a worst-case dominating pair if additionally $P=\mathcal{M}(D),Q=\mathcal{M}(D^{\prime})$ for some neighbouring datasets $D,D^{\prime}$.
In practice, a typical privacy accounting is done by taking each $L_{i}$ to be the PLD of a worst-case dominating pair of the mechanism $\mathcal{M}_{i}$. We observe below that such PLD is symmetric. Thus, the assumption in Section˜7 is quite a natural one that is already satisfied in practice.
Lemma 15.
If $(P,Q)$ is a worst-case dominating pair of $\mathcal{M}$, then $(P,Q)$ is symmetric.
@@ Appendix C On Symmetric PLDs and Dominating Pairs
Since $(P,Q)$ is a worst-case dominating pair, we have $P=\mathcal{M}(D)$ and $Q=\mathcal{M}(D^{\prime})$ for some neighbouring datasets $D,D^{\prime}$ and $(P,Q)\succeq(\mathcal{M}(D^{\prime}),\mathcal{M}(D))=(Q,P)$. By definition, there exists a Markov kernel $\kappa$ such that $\kappa P=Q$ and $\kappa Q=P$, but this also implies that $(Q,P)\succeq(P,Q)$. ∎
We will next prove Lemma˜8. To do so, let us start by stating a relationship between the hockey-stick divergence at $x$ and $x^{-1}$, which will be convenient for our proof.
Lemma 16 ([58]).
For any pair $(P,Q)$ and $x\in\mathbb{R}^{\times}$, we have $H_{x}(Q\parallel P)=x\cdot H_{x^{-1}}(P\parallel Q)-x+1$.
Proof of Lemma˜8.
Let $L=\operatorname{PLD}(P,Q),L^{\prime}=\operatorname{PLD}(P^{\prime},Q^{\prime})$ where $(P,Q),(P^{\prime},Q^{\prime})$ are symmetric pairs. The forward direction is obvious (from definition of dominating). For the backward direction, if $0<x<1$, then Lemma˜16 implies that
$H_{x}(P\parallel Q)\geq H_{x}(P^{\prime}\parallel Q^{\prime})\Leftrightarrow H_{x^{-1}}(Q\parallel P)\geq H_{x^{-1}}(Q^{\prime}\parallel P^{\prime})\Leftrightarrow H_{x^{-1}}(P\parallel Q)\geq H_{x^{-1}}(P^{\prime}\parallel Q^{\prime}),$
where the second equivalence is due to the symmetry assumption. ∎